# Reservoir-Computing Model for Mapping and Forecasting Neuronal Interactions from Electrophysiological Data

### 23 Jan 2024 v2

@@ 4.4 Linearized Model and Connectivity Analysis
Let us observe the dynamics of the model. If the initial state of the reservoir is $\mathbf{x}[0]=0$ (unexcited state), we note that, without any input $\mathbf{y}$, the time sequence of the reservoir state, $\mathbf{x}[n]$, (2b), will not change its state and as a consequence, according to (2.1), no dynamics in the network nodes $\mathbf{y}[n]$ will be observed. Let us assume (without the loss of generality) that at a certain time-step $n=1$ we have a small perturbation, $\mathbf{y}[1]$, such that according to (2b) the following holds:
@@
This approximation is valid for any nonlinear function that satisfy $f(\xi)\approx\xi$ for $\xi\ll 1$, e.g., (14). If such linear regime is maintained in the following $k-1$ steps, from (2.1) it yields that the network output at time step $k+1$ is:
@@
is a $N_{ch}\times N_{ch}$ transfer matrix of order $p$ (note that we omitted the constant bias vector $\mathbf{b}$ from (2c), since it describes a constant DC offset, and as known a posteriori, its value is small).
Assuming that the regression (Eqs. (2.1),(4)) as part of the model training has achieved low training and validation error score, means that a feasible parametrization for the equations of the nonlinear model (2.1) has been found. It follows that the transfer matrices (22) contain the connectivity weights between the nodes in the linear regime, for different orders of interaction, where their effect scales as $\alpha^{p}$
Note that by eliminating the reservoir operation, i.e., canceling the effect of previous steps, that is, taking $\alpha\to 0$ in (21), will lead to the following equation:
$\mathbf{y}[k+1]=\mathcal{T}_{0}\mathbf{y}[k]$ (23)
from which we define the intrinsic connectivity matrix $\mathcal{T}_{0}$ described by Eq. (6), since it describes directly the weights between the network nodes for two consecutive states, regardless of the memory stored in the reservoir. Note that each component $\mathcal{T}^{i,j}_{0}$ shows the directed connection $j\rightarrow i$, i.e., from node $j$ at time $n$ to node $i$ at time $n+1$. The higher order $\mathcal{T}_{p}$ ($p=1,2,\ldots k$) matrices contain the corrections (still in the linear regime) to the connection weights following the reservoir activation. These transfer matrices express both excitatory connections (positive values) and inhibitory connections (negative values).
@@ 4.5 Testing and Performance Metrics
Model performance was examined by several validation paths: 1. Evaluating training (Eq. (2.1)) and validation loss (Eq. (4.3.2) while tuning model parameters $\alpha$, $m$ and the integration time (see Supplementary Materials). 2. Benchmarking the connectivity matrix using synthetic data; 3. Estimating the prediction accuracy of the network spatio-temporal response using experimental and synthetic data.
@@ 4.5.1 Connectivity Map
By obtaining the Intrinsic Connectivity Matrix, $\mathcal{T}_{0}$ from Eq. (6), we are able to build a network graph corresponding to the interactions between the nodes of the given network. We assume that the intrinsic connections, described by the weights of $\mathcal{T}_{0}$, depict the main component of the short term interaction between the nodes, and hence they describe, to some extent, the effective connectivity between the different populations (or circuits) in the network.
To assess the accuracy of the connectivity predictions derived from Eq. (6), we utilized two categories of metrics: 1. binary metrics (determining the presence or absence of connections) and 2. weight prediction metrics. The former was assessed using ROC curves, while the latter was evaluated through Pearson Correlation Coefficient, $\rho(X,Y)$. Ground truth connections were established as an input to NEST simulation, as detailed in Section 4.6. This simulation generated electrophysiological data based on the provided structural information, which was subsequently employed for training the model in order to derive the connectivity map. It’s important to emphasize that the comparison of connectivity can only be conducted using synthetic data, as this specific information cannot be derived a priori from a biological neuronal culture without additional processing.
In the ROC analysis, we designated ’1’ for any non-zero weights in the ground truth connectivity matrix, denoted as $\mathcal{T}_{GT}$, and ’0’ for zero values. In contrast, the model-predicted connections were categorized as ’1’ or ’0’ using a range of custom thresholds based on the evaluation of True Positive Rate (TPR) versus False Positive Rate (FPR) for each of these thresholds.
As for the more precise prediction metric, we used the Pearson Correlation Coefficient, since it evaluates the similarity of two multi-dimensional datasets regardless of the scale of each one of them. We therefore calculated $\rho(\mathcal{T}_{0},\mathcal{T}_{GT})$, where $\mathcal{T}_{0},\mathcal{T}_{GT}$ are the intrinsic connectivity matrix obtained by the model (Eq. (6)) and the ground truth matrix, respectively. Note that for the propriety of calculation of $\rho$, $\mathcal{T}_{0}$ and $\mathcal{T}_{GT}$ are transformed into 1D vectors, where each component corresponds to a weight of a specific connection.
We benchmarked our model against connectivity matrices obtained from previously developed methods. Notably, the Cross-Correlation (CC) and Transfer Entropy (TE) methods emerged as the most effective approaches among the existing ones. To obtain the connectivity matrices generated by these methods, denoted as $\mathcal{T}_{CC}$ and $\mathcal{T}_{TE}$ respectively, we employed the SpiCoDyn toolbox, as described in Reference [34]. Subsequently, we estimated their ROC curve against the ground truth connectivity and calculated the correlation coefficients $\rho(\mathcal{T}_{CC},\mathcal{T}_{GT})$ and $\rho(\mathcal{T}_{TE},\mathcal{T}_{GT})$ to measure their similarity against the ground truth connectivity matrix.
@@ 4.5.2 Response Prediction Test
As outlined in Section 2.3, our approach for this test involved a distinctive procedural path. In this part, we implemented a methodology wherein the training and validation phases were executed on a dataset comprising spontaneous activity of a culture. Following this, the trained model underwent a response test, scrutinizing its reaction to a specific input and observing its temporal propagation, as illustrated by Eq. (8). Finally, the model’s response was evaluated based on its agreement with either experimental or simulated data characterizing the response of the neuronal network to an external impulse corresponding to the modeled input. It is worth noting that the modeled input is not characterized in a fully physical form, such as a pulse with a defined temporal profile, rather it is represented as an input vector with dimensions $N_{ch}\times 1$, translated to the model attributes, where the spatial elements of the vector correspond to the spatial representation of the macroscopic network (the electrodes layout); the one column representation infers that the impulse is given during one time-step persisting for one integration time; and the amplitude of the stimulus corresponds to the normalized instantaneous spike-rate, as was coded in the training data. This representation of the impulse describes the effective pre-synaptic impulse given around the target population.
As the test data, we prepared datasets from experiments of evoked activity driven by a localized stimulus (optical, electrical or simulated on NEST). The experimental response is a recorded time trace of the network activity within a fixed time window following a stimulus, presenting the ISR in small time bins, corresponding to one time step set as the integration time of the model, and averaged over numerous stimuli. This time histogram is known in literature as Post-Stimulus Time-Histogram (PSTH).
Also here, as in the connectivity prediction, we employ a few types of metrics in order to evaluate the prediction of the modeled network response:
@@
    ROC Curve: used to assess the spatial prediction accuracy of responses by utilizing binary classification, distinguishing between responsive channels/electrodes and unresponsive ones. In establishing the ground truth for the ROC curve, we designated a channel as ’1’ (indicating a responsive electrode) if it displayed, in at least half of the repeated stimuli, at least one time bin in the PSTH with a value corresponding to one spike per this time bin; and ’0’ otherwise. The prediction values were determined as the maximum value among the elements of the matrix $\tilde{\mathbf{Y}}$, representing the peak of each channel’s predicted PSTH. Subsequently, the ROC curve analysis was conducted in a manner consistent with the procedure described above for the connectivity map prediction metric.
@@
    Cross-Root-Mean-Squared-Error (XRMSE): used to evaluate the spatio-temporal prediction accuracy by examining the temporal profile of the response for each channel/electrode. Due to the hypothesized disparities described above between the physical and the modeled input signals, we expected the modeled response to have certain inaccuracies in the temporal profile of each single electrode, such as temporal distortion of the signal or time shift. To take into account these considerations, we followed the following accuracy evaluation procedure. Let $\tilde{\mathbf{U}}=\bigl{[}\tilde{\mathbf{u}}[1],\tilde{\mathbf{u}}[2],\ldots,\tilde{\mathbf{u}}[N_{t}]\bigr{]}$ and $\mathbf{U}=\bigl{[}\mathbf{u}[1],\mathbf{u}[2],\ldots,\mathbf{u}[N_{t}]\bigr{]}$ be the predicted and the experimental PSTH, respectively, describing a response to an input $\mathbf{u}_{in}$. It is important to note that the dataset $\mathbf{U}$ is normalized by the same normalization factor of the basal activity of the corresponding network (Sec. 4.3.2). We then perform the following calculations:
    • The integrated response for each electrode (approximated by trapezoid integration):
@@
      where $U_{i,n}$ represents the element $i,n$ of the matrix $\mathbf{U}$; first element is the channel index and second element is the time step index.
    • Weighting each channel by its response:
@@
      where $(\varsigma_{i},\tilde{\varsigma}_{i})$ are the integrated response for the observed and the model predicted time trace at channel $i$, respectively.
    For the evaluating the accuracy of the predicted response, we use the following metrics:
    • For each pair $(\mathbf{U},\tilde{\mathbf{U}})$ we define the Cross-root-mean-squared error function (XRMSE) by:
      $XRMSE_{i}[\tau]=\sqrt{\sum_{n=1}^{N_{t}}\upsilon_{i,n,\tau}\left(U_{i,n}-\tilde{U}_{i,n-\tau}\right)^{2}}$ (26)
      where $\tau\in[-\tau_{max},+\tau_{max}]$ is the time shift between the observation time trace $U$ and the predicted time trace $\tilde{U}$, $\tau_{max}$ is the maximal time shift (in time steps) taken for estimation (in our calculations we set $\tau_{max}=10$); $\upsilon_{i,n,\tau}$ is a weight coefficient of each time point, expressed by:
      $\upsilon_{i,n,\tau}=\frac{U_{i,n}+\tilde{U}_{i,n-\tau}}{\sum_{n}\left(U_{i,n}+\tilde{U}_{i,n-\tau}\right)}$ (27)
    • Finding the time lag in which the error is minimal:
[…]
    • Evaluation of the aggregated error of the response prediction:
      $\bar{R}=\sum_{i}\chi_{i}\cdot\varepsilon_{i}$ (29)
      which leads to Eq. (9); and the aggregated time lag of the prediction:
      $\bar{\tau_{l}}=\sum_{i}\chi_{i}\cdot\tau_{l,i}$ (30)
Below are the considerations for the $\bar{R}$ metric:
  • The XRMSE metric was introduced to assess the predictive accuracy of signals featuring a peak and to account for temporal misalignment.
  • The introduction of weights $\upsilon_{i,n,\tau}$ serves the purpose of diminishing the significance of zero or low values, as recurrent predictions in subsequent points are considered trivial.
  • The introduction of the weights $\chi_{i}$ was aimed at assigning greater significance to channels exhibiting stronger responses.
@@ 4.6 In-silico Simulation
As ground truth connections among neuronal populations are not available for experimental recordings of neuronal electrophysiological signals (such as MEA recording), an in-silico model of neuronal dynamics has been developed, thus allowing evaluation of the RCN prediction. The basic network unit of such model is defined as population, which is composed of a fixed number of point-process neurons described by Izhikevich equations [35], parameterized to follow the dynamic of regular-spiking neurons [36]. Each neuron inside a population is connected to all neighboring neurons (intra-population connections), with no self-connections allowed, and the different populations are then connected randomly to a variable number of the others (inter-population connections). Specifically, a connection between two populations $i$ and $j$ is obtained by connecting each neuron of population $i$ to $n$ randomly sampled neurons of population $j$. This network architecture has been chosen to recreate highly interconnected hubs[37], represented by single populations, which should resemble neuronal assembles surrounding actual MEA electrodes for which action potentials are recorded. Instead, connections among different populations mimic long-range relationships between neurons. For both intra- and inter-population connections, different types of synapses — static and plastic [38] — have been tested. The weights are assigned using uniform distributions, with distinct excitatory and inhibitory connection ranges. In particular, inhibitory connection weights have a higher absolute value to accommodate their lower prevalence in the network, maintaining an 80/20% proportion between excitatory and inhibitory connections. To replicate the spontaneous activity displayed by neuronal cultures, a piecewise constant current with a Gaussian-distributed amplitude has been injected into each neuron to reproduce the fluctuation in the membrane potential (noise component). Moreover, Poisson spike trains addressed either to all neurons or just one neuron inside populations, have been utilized to replicate spikes occurring from neurons in the culture not detected by MEA electrodes (background activity). The model has been tuned to reproduce the dynamics exhibited by in-vitro neuronal networks. In particular, simulations show a mix of spiking and bursting activities as visible by the raster plot depicted in Fig.8(a), with an average firing rate (AFR - spikes/s) in line with the experimental recordings. Moreover, the log-inter-spike-interval distribution (log-ISI) has been calculated for each population, as visible in Fig.8(b), to monitor the dynamic of the network, as in general in-vitro cultures showing bursting activity are characterized by electrodes whose log-ISI histograms appear bimodal [39, 31]. The final output of this simulation is obtained by collecting spike times of each population’s neuron and by sorting them temporally, thus obtaining a spike train for each population, which resembles the spike train acquired after performing spike detection on the raw multi-unit activity signal sampled by a MEA electrode[40]. Additionally, a connectivity matrix (CM) outlining links between populations is obtained. In particular, each entry $i,j$ indicates a link between populations $i$ and $j$, and its weight value is a weighted average of all inter-population connection weights by the firing probability of every source neuron in population $i$. The weighted average has been selected to take into account not just structural connections but also the functional relationships across populations. Moreover, to test the ability of the RCN to retrieve the connectivity of the in-silico model, networks composed of clusters of populations have been created, where inter-population connections were allowed only between populations belonging to the same cluster. Furthermore, populations whose inter-population connections were only inhibitory have been used to evaluate the capacity of the RCN to distinguish between excitatory and inhibitory synapses.
Figure 9: (a) Raster Plot of simulated electrophysiological activity of an in-silico network composed of 16 populations. (b) example of log-ISI distribution of population 0 of the same simulation depicted in the raster plot. The log-ISI histogram has been obtained as described in [31].
@@
Moreover, as the RCN can be used to predict electrodes’ responses to a specific stimulus, we used the in-silico model to simulate such responses. For each stimulation protocol, a constant DC input lasting 2 ms was provided to all neurons of a specific population, for a total of 10 equally-spaced repetitions across 10 seconds length. One, or more stimulation protocols have been performed for each simulation by selecting a different targetpopulation, with 1 second of separation between each stimulation protocol. The starting and ending times of each repetition have been later used to compute the PSTH of the stimulated population. Different configurations have been tested with networks composed from 4 up to 60 populations. Each simulation was generated and simulated with different seed values for the random number generator of our simulation, leading each time to a different set of both intra- and inter-population connections. Moreover, each simulation lasted for five minutes, divided into two sections: a "background activity" section, lasting half of the recording, where only the noise and background components were active, and the "stimulation" section, where stimulation protocols could be performed in addition to the basal components. In particular, the background activity data was later used to train the RCN model and to calculate the Intrinsic Connectivity Matrix, while the stimulation data to perform response prediction tests.
Statistic
@@
For Figure 3 (b), a Kruskal-Wallis H-test has been performed to calculate if the population median of all of the groups are equal considering each metric. Subsequently, a series of two-sided Mann-Whitney-Wilcoxon test with Bonferroni correction have been used to compare pairwise the algorithms performance.
@@
All simulations were performed using the NEST simulator (version 3.5) [41] with custom code implemented in Python (version 3.8.17), with a kernel resolution of 0,01 ms. RCN model has been developed in Python implementing new classes on top of the ReservoirPy library (version 0.3.9) [42]. Common Python library, such as Numpy, Scipy, and Scikit-learn have been used throughout the custom code for analysis.
@@
The code for simulations, RCN model, and data analyses in this work will be available in public repository. Code can be shared upon request.
@@
The data of the simulations, recordings and data analyses reported in this study will be available in a public repository. Data can be shared upon request.

### 05 Mar 2024 v3

@@ 5.4 Linearized Model and Connectivity Analysis
If the initial state of the reservoir is $\mathbf{x}[0]=0$ (resting state) and in the absence of any input $\mathbf{y}$, the reservoir state $\mathbf{x}[n]$ does not change with time, see Eq. (2b). As a consequence, no dynamics in the network nodes $\mathbf{y}[n]$ is observed, see Eq. (2). Let us assume (without the loss of generality) that at a certain time-step $n=1$ we have a small perturbation, $\mathbf{y}[1]$, such that:
@@
This approximation is valid for any nonlinear function that satisfies $f(\xi)\approx\xi$ for $\xi\ll 1$, e.g., Eq. (14). If such linear regime is maintained in the following $k-1$ steps, from Eq. (2) follows that the network output at time step $k+1$ is:
@@
is a $N_{ch}\times N_{ch}$ transfer matrix of order $p$. Note that we omitted the constant bias vector $\mathbf{b}$ from Eq. (2c), since it describes a constant DC offset, and as it results a posteriori, its value is negligibly small.
After training the RC model network, a proper parametrization of the nonlinear model given by Eq. (2) is found. It follows that the $p$ transfer matrices (22) contain the connection strengths (elements of the matrices) between the nodes for different $p$ steps. These interactions scale with time according to $\alpha^{p}$. Next, if we eliminate the recurrence in the reservoir operation by removing the effect of previous steps (no memory in the reservoir, i.e. $\alpha\to 0$ in Eq. (21)), we get:
$\mathbf{y}[k+1]=\mathcal{T}_{0}\mathbf{y}[k]\;.$ (23)
We define $\mathcal{T}_{0}$ the intrinsic connectivity matrix since it directly gives the weights between the network nodes for two consecutive temporal states, regardless of the memory stored in the reservoir. Indeed, each matrix element $\mathcal{T}^{i,j}_{0}$ shows the direct connection $j\rightarrow i$, i.e., from node $j$ at time $n$ to node $i$ at time $n+1$. The higher order $\mathcal{T}_{p}$ ($p=1,2,\ldots k$) matrices contain the corrections to the connection weights following the reservoir activation still in the linear approximation. The elements of $\mathcal{T}_{0}$ express both excitatory (positive values) as well as inhibitory connections (negative values).
@@ 5.5 Testing and Performance Metrics
Model performance was examined by several validation paths: 1. Evaluating training (Eq. (2)) and validation loss (Eq. (5.3.2) while tuning model parameters $\alpha$, $m$ and the integration time (see Supplementary Materials). 2. Benchmarking the connectivity matrix using synthetic data; 3. Estimating the prediction accuracy of the network spatio-temporal response to a localized stimulus using experimental and synthetic data.
@@ 5.5.1 Connectivity Map
By obtaining the Intrinsic Connectivity Matrix, $\mathcal{T}_{0}$ from Eq. (7), we build a network graph that describes the interactions between the network nodes. We assume that these connections depict the short-term interaction between the nodes, and, hence, the effective connectivity between the different populations (or circuits) in the neuronal culture. To assess the accuracy of the network graph, we utilized two metrics: i) a binary metric, i. e. the presence or absence of connections, assessed by ROC curves, and ii) a weight prediction metric, evaluated through the Pearson Correlation Coefficient, $\rho(X,Y)$. In the ROC analysis, we designated ’1’ for any non-zero weights in the ground-truth network connectivity matrix, denoted as $\mathcal{T}_{GT}$, and ’0’ for zero values. In contrast, the connections for the RC model were categorized as ’1’ or ’0’ using a range of custom thresholds based on the evaluation of True Positive Rate (TPR) versus False Positive Rate (FPR). As for the Pearson Correlation Coefficient, we calculated $\rho(\mathcal{T}_{0},\mathcal{T}_{GT})$, where $\rho$, $\mathcal{T}_{0}$ and $\mathcal{T}_{GT}$ are transformed into 1D vectors of weights for a specific connection. We benchmarked the RC model against connectivity matrices obtained from the Cross-Correlation (CC) and Transfer Entropy (TE) methods. To obtain the connectivity matrices generated by these methods, denoted as $\mathcal{T}_{CC}$ and $\mathcal{T}_{TE}$ respectively, we employed the SpiCoDyn toolbox, as described in [39]. Subsequently, we estimated their ROC curve against the ground-truth network connectivity and calculated the correlation coefficients $\rho(\mathcal{T}_{CC},\mathcal{T}_{GT})$ and $\rho(\mathcal{T}_{TE},\mathcal{T}_{GT})$.
@@ 5.5.2 Response Prediction Test
As outlined in Section 3.2, a trained model undergoes a response test where the stimulus is represented as an input vector with dimensions $N_{ch}\times 1$. The elements of the vector correspond to the spatial representation of the macroscopic domain (the electrodes layout); the one-column representation infers that the impulse is given during one time-step persisting for one integration time; and the amplitude of the stimulus corresponds to the normalized instantaneous spike-rate, as was coded in the training data. This representation of the impulse describes the effective pre-synaptic impulse given to a target neuronal population that insists on the electrode-probed region. As the test data, we prepared datasets from experiments of evoked activities driven by a localized stimulus (optical, electrical, or simulated on NEST). The experimental response is a recorded time trace of the neuronal culture activity within a fixed time window following the stimulus in the form of a time histogram, which is known in the literature as post-stimulus time-histogram (PSTH).
The following metrics evaluate the accuracy of the prediction of the RC model network’s response:
@@
    ROC Curve: used to assess the accuracy of spatial responses by utilizing a binary classification. This distinguishes between responsive channels/electrodes and unresponsive (resting) ones. In establishing the experiment’s ROC curve, we designated a channel as ’1’ (indicating a responsive electrode) if at least one time bin in the PSTH has a value corresponding to one spike per time bin in at least half of the repeated stimuli; and ’0’ otherwise. The predicted values were determined as the maximum value among the elements of the matrix $\tilde{\mathbf{Y}}$, representing the peak of each node’s predicted PSTH. Subsequently, the ROC curve analysis was conducted as for the connectivity map analysis.
@@
    An aggregated error $\bar{R}$ based on the root-mean-squared-error: used to evaluate the spatio-temporal prediction accuracy by examining the temporal profile of the response for each channel/electrode or nodes.
Due to the differences between the experimental and the modeled stimuli, we expected inaccuracies in the temporal profile of the node response, such as temporal distortion of the signal or time shift. To take into account this, we used the following procedure to compute $\bar{R}$. Let $\tilde{\mathbf{U}}=\bigl{[}\tilde{\mathbf{u}}[1],\tilde{\mathbf{u}}[2],\ldots,\tilde{\mathbf{u}}[N_{t}]\bigr{]}$ and $\mathbf{U}=\bigl{[}\mathbf{u}[1],\mathbf{u}[2],\ldots,\mathbf{u}[N_{t}]\bigr{]}$ be the predicted and the experimental PSTH, respectively. $\mathbf{U}$ is normalized by the same normalization factor as the basal activity for a given culture (Sec. 5.3.2). We then compute the integrated response for each electrode (approximated by a trapezoid integration):
@@
where $U_{i,n}$ represents the $i,n$ elements of the matrix $\mathbf{U}$: the first index is the electrode index and the second index is the time step index. Next, we define a weight for each node/channel, as follows:
@@
where $(\varsigma_{i},\tilde{\varsigma}_{i})$ are the integrated response for the experimental and the model time traces, respectively. Then, we define $R_{i}$ the cross root-mean-square- error per channel, which can be calculated by:
$R_{i}[\tau]=\sqrt{\sum_{n=1}^{N_{t}}\upsilon_{i,n,\tau}\left(U_{i,n}-\tilde{U}_{i,n-\tau}\right)^{2}}$ (26)
where $\tau\in[-\tau_{max},+\tau_{max}]$ is the time shift between the experimental time trace $U$ and the predicted time trace $\tilde{U}$, $\tau_{max}=10$ is arbitrarily used; $\upsilon_{i,n,\tau}$ is a temporal weight coefficient given by:
$\upsilon_{i,n,\tau}=\frac{U_{i,n}+\tilde{U}_{i,n-\tau}}{\sum_{n}\left(U_{i,n}+\tilde{U}_{i,n-\tau}\right)}\;.$ (27)
Finally, by finding the time lag in which the error is minimal:
$\begin{gathered}\varepsilon_{i}=\min(R_{i}[\tau])\\ \tau_{l,i}={\arg\min}(R_{i}[\tau])\;,\end{gathered}$ (28)
we can evaluate the aggregated error $\bar{R}$ :
$\bar{R}=\sum_{i}\chi_{i}\cdot\varepsilon_{i}\;,$ (29)
which is Eq. (9), and the aggregated time lag of the prediction:
$\bar{\tau_{l}}=\sum_{i}\chi_{i}\cdot\tau_{l,i}\;.$ (30)
In the evaluation of $\bar{R}$, we used Eq. (26) to assess the predictive accuracy of signals featuring a peak and to account for temporal misalignment; the weights $\upsilon_{i,n,\tau}$ serve to diminish the significance of zero or low values, as recurrent predictions in subsequent points are considered trivial; the weights $\chi_{i}$ assign a greater significance to channels exhibiting stronger responses.
@@ 5.6 In-silico Simulation
As the underlying circuital network of a neuronal culture is not accessed by the experimental recordings of the neuronal electrophysiological signals (MEA recording), an in-silico model of a neuronal culture has been developed. The basic network unit of such model is defined as population, which is composed of a fixed number of point-process neurons described by Izhikevich equations [40], parameterized to follow the dynamic of regular-spiking neurons [41]. Each neuron inside a population is connected to all neighboring neurons (intra-population connections), with no self-connections allowed. The different populations are then connected randomly to a variable number of the others (inter-population connections). Specifically, a connection between two populations $i$ and $j$ is obtained by connecting each neuron of population $i$ to $n$ randomly sampled neurons of population $j$. This network architecture has been chosen to recreate highly interconnected hubs[42], represented by single populations, which should resemble neuronal assemblies surrounding actual MEA electrodes for which action potentials are recorded. Instead, connections among different populations mimic long-range relationships between neurons. For both intra- and inter-population connections, different types of synapses — static and plastic [43] — have been tested. The weights are assigned using uniform distributions, with distinct excitatory and inhibitory connection ranges. In particular, inhibitory connection weights have a higher absolute value to accommodate their lower prevalence in the network, maintaining an 80/20% proportion between excitatory and inhibitory connections. To replicate the spontaneous activity displayed by neuronal cultures, a piecewise constant current with a Gaussian-distributed amplitude has been injected into each neuron to reproduce the fluctuation in the membrane potential (noise component). Moreover, Poisson spike trains addressed either to all neurons or just one neuron inside populations, have been utilized to replicate spikes occurring from neurons in the culture not detected by MEA electrodes (background activity). The model has been tuned to reproduce the dynamics exhibited by in-vitro neuronal networks. In particular, simulations show a mix of spiking and bursting activities as visible by the raster plot depicted in Fig.4(a), with an average firing rate in line with the experimental recordings. Moreover, the log-inter-spike-interval distribution (log-ISI) has been calculated for each population to monitor the network’s dynamic. Populations showed different firing dynamics, resulting in heterogenous log-ISI histograms, with the appearance of bimodal histograms (Fig. 4(b)), as observed in-vitro neuronal cultures [44, 36], generally indicating bursting events. The final output of in-silico simulation is obtained by collecting spike times of each neuron in a population and by sorting them temporally, thus obtaining a spike train for each population, which resembles the spike train acquired after performing spike detection on the raw multi-unit activity signal sampled by a MEA electrode [45]. Additionally, the connectivity matrix of the links between populations can be obtained. In particular, each entry $i,j$ indicates a link between populations $i$ and $j$, and its weight value is a weighted average of all inter-population connection weights by the firing probability of every source neuron in population $i$. The weighted average has been selected to take into account not just structural connections but also the functional relationships across populations. Moreover, to test the ability of the RC model to retrieve the connectivity of the in-silico model, networks composed of clusters of populations have been created, where inter-population connections were allowed only between populations belonging to the same cluster. Furthermore, populations whose inter-population connections were only inhibitory have been used to evaluate the capacity of the RC model to distinguish between excitatory and inhibitory synapses.
@@
In addition, on the in-silico model localized stimulation simulations have been performed. For each stimulation protocol, a constant DC input lasting 2 ms was provided to all neurons of a specific population, for a total of 10 equally-spaced repetitions across 10 seconds length. One, or more stimulation protocols have been performed for each simulation by selecting a different targetpopulation, with 1 second of separation between each stimulation protocol. Each repetition’s starting and ending times have been later used to compute the PSTH of the stimulated population. Different configurations have been tested with networks of 4 up to 60 populations. Each simulation was generated and simulated with different seed values for the random number generator of our simulation, leading each time to a different set of both intra- and inter-population connections. Moreover, each simulation lasted for five minutes, divided into two sections: a background activity section, lasting half of the recording, where only the noise and background components were active, and the stimulation section, where stimulation protocols could be performed in addition to the basal components. In particular, the background activity data was later used to train the RC model and to calculate the Intrinsic Connectivity Matrix, while the stimulation data to perform response prediction tests.
Statistics
@@
For Fig. 3(b), a Kruskal-Wallis H-test has been performed to calculate if the population median of all groups is equal considering each metric. Subsequently, a series of two-sided Mann-Whitney-Wilcoxon tests with Bonferroni correction has been used to compare pairwise the algorithms’ performance.
@@
All simulations were performed using the NEST simulator (version 3.5) [14] with custom code implemented in Python (version 3.8.17), with a kernel resolution of 0.01 ms. The RC model has been developed in Python implementing new classes on top of the ReservoirPy library (version 0.3.9) [46]. Common Python libraries, such as Numpy, Scipy, and Scikit-learn have been used throughout the custom code for analysis. SpiCoDyn version 3.3 [39] has been used to perform Cross-Correlogram (CC) and Transfer Entropy (TE) analysis.
@@
The code for simulations, the RC model, and data analyses in this work will be available in a public repository. The code will be shared upon request.
@@
The data of the simulations, recordings, and data analyses reported in this study will be available in a public repository. Data can be shared upon request.